OpenStudy (anonymous):

whats the equation of a line that passes through (0,1) and (2,3)

OpenStudy (akashdeepdeb):

If you have 2 poitns you can find the slope right! RISE AND RUN? (y2-y1)/(x2-x1) FIND THAT SLOPE(m) And then use equation (y-y1) = m(x-x1) WHERE (x1,y1) = any of your points!! :) I hope you get it!